# Annealed Multiple Choice Learning: Overcoming limitations of Winner-takes-all with annealing

### v1

@@ D.3.3 Phase transition in aMCL training
where $\rho=0.9$ is the decay factor, and $T_{0}\in[0.1,5,23]$ is the initial temperature. Figure 6 indicates the negative MCL SI-SDR score computed during training on a validation set. In order to compare meaningfully the different temperature schedules, we plot the separation score against the temperature, regardless of the training step when this temperature is reached.
We see that that aMCL exhibits phase transitions in this setting. Indeed, when we use a scheduler with high initial temperature $T_{0}\in[5,23]$, the negative MCL SI-SDR metric stays on a plateau and then decreases after some critical temperature $T_{0}^{c}$ has been reached. As expected, this critical temperature seems to be independent on the initial temperature $T_{0}>T_{0}^{c}$. Moreover, when using a scheduler with initial temperature $T_{0}\approx 0.1<T_{0}^{c}$, aMCL does not undergo a phase transition. This suggests that the phase transition phenomenon exhibited in Section A also exists in the non-Euclidean case, even though the properties of the barycenter and the shape of the Voronoi cells have to be redefined. Further work will analyze more exhaustively the impact of the scheduler type on this phase transition phenomenon. We hypothesize that the temperature decay speed at the critical temperature $T_{0}^{c}$ will play a key role in this analysis.

### v3

@@ E.3.3 Phase transition in aMCL training
where $\rho=0.9$ is the decay factor, and $T_{0}\in\{0.1,5,23\}$ is the initial temperature. Figure 8 indicates the negative MCL SI-SDR score computed during training on a validation set. In order to compare meaningfully the different temperature schedules, we plot the separation score against the temperature, regardless of the training step when this temperature is reached.
We see that that aMCL exhibits phase transitions in this setting. Indeed, when we use a scheduler with high initial temperature $T_{0}\in\{5,23\}$, the negative MCL SI-SDR metric stays on a plateau and then decreases after some critical temperature $T_{0}^{c}$ has been reached. As expected, this critical temperature seems to be independent of the initial temperature $T_{0}>T_{0}^{c}$. Moreover, when using a scheduler with initial temperature $T_{0}\approx 0.1<T_{0}^{c}$, aMCL does not undergo a phase transition. This suggests that the phase transition phenomenon exhibited in Section A also exists in the non-Euclidean case, even though the properties of the barycenter and the shape of the Voronoi cells have to be redefined. Further work will analyze more exhaustively the impact of the scheduler type on this phase transition phenomenon. We hypothesize that the temperature decay speed at the critical temperature $T_{0}^{c}$ will play a key role in this analysis.